OpenStudy (anonymous):

Under certain conditions, neon (Ne) gas diffuses at a rate of 7.0 centimeters per second. Under the same conditions, an unknown gas diffuses at a rate of 4.9 centimeters per second. What is the approximate molar mass of the unknown gas?

OpenStudy (aaronq):

Do you know Graham's Law of diffusion?

OpenStudy (anonymous):

That I do not know..

OpenStudy (aaronq):

it this:\[\frac{ Rate _{1} }{ Rate _{2} }=\sqrt{\frac{ M _{2} }{M _{1} }}\] M=molar mass and it works because the rate of diffusion is proportional to the mass of the molecules.

OpenStudy (anonymous):

Okay lemme try

OpenStudy (aaronq):

aight

OpenStudy (anonymous):

41.2 g/mol is what I got

OpenStudy (aaronq):

it's good, now look at the periodic table for a gas that would of similar molar mass

OpenStudy (anonymous):

I'm feeling a dose of Molybdenum

OpenStudy (aaronq):

no :P molybdenum is like 95 g/mol

OpenStudy (anonymous):

Aight so calcium is the next best thing

OpenStudy (anonymous):

It's the closest match I could find

OpenStudy (aaronq):

yeah, i guess that would be the closest thing. I just re-read the question and i thought you had to identify the gas lol but you don't.

OpenStudy (anonymous):

Alright so I have a question that I already answered but I want to make sure it's correct

OpenStudy (anonymous):

If 2.0 liters of oxygen gas react with excess carbon monoxide at STP, how many liters of carbon dioxide can be produced under the same conditions?

OpenStudy (anonymous):

It's multiple choice and the answer I got was 4.0L

OpenStudy (aaronq):

yeah, thats right.

OpenStudy (anonymous):

Worddd thank you

OpenStudy (aaronq):

no probs !